three students select cards from a standard deck, without replacement. find the probability that: a)all the cards have the same suit b) all the cards are of differnt suits
Answer by stanbon(75887)

three students select cards from a standard deck, without replacement. find the probability that:
-------
a) all the cards have the same suit
# of ways to succeed: 4*4C3 = 4*4 = 16
# of random selections: 52C3
----
P(3 card of same suit) = 16/52C3
----------------------------------------
b) all the cards are of different suits
(1/52)*(1/39)*(1/26) = 1/52728
========================================
Cheers,
Stan H.